OpenStudy (anonymous):

multiply using "FOIL" ? (t – 1/5)^2

OpenStudy (anonymous):

The ^2 means square, which means to multiply the things by itself. So, that means, (t- (1/5)) * (t- (1/5))

OpenStudy (anonymous):

First Inner Outer Last (t-1/5)*(t-1/5) = t*t +-1/5t -1/5t +(-1/5)(-1/5) = t^2 -2/5t +1/25

OpenStudy (anonymous):

thank you!